# When Less Language is More: Language-Reasoning Disentanglement Makes LLMs Better Multilingual Reasoners

### v1

@@ Abstract
Multilingual reasoning remains a significant challenge for large language models (LLMs), with performance disproportionately favoring high-resource languages. Drawing inspiration from cognitive neuroscience, which suggests that human reasoning functions largely independently of language processing, we hypothesize that LLMs similarly encode reasoning and language as separable components that can be disentangled to enhance multilingual reasoning. To evaluate this, we perform a causal intervention by ablating language-specific representations at inference time. Experiments on 10 open-source LLMs spanning 11 typologically diverse languages show that this language-specific ablation consistently boosts multilingual reasoning performance. Layer-wise analyses further confirm that language and reasoning representations can be effectively decoupled throughout the model, yielding improved multilingual reasoning capabilities, while preserving top-layer language features remains essential for maintaining linguistic fidelity. Compared to post-training such as supervised fine-tuning or reinforcement learning, our training-free ablation achieves comparable or superior results with minimal computational overhead. These findings shed light on the internal mechanisms underlying multilingual reasoning in LLMs and suggest a lightweight and interpretable strategy for improving cross-lingual generalization. Refer to caption Figure 1: Overview of our hypothesis and empirical findings. Left: Motivated by cognitive neuroscience, we hypothesize that reasoning and language processing in LLMs can be disentangled. Right: To validate this hypothesis, we perform a causal intervention by removing language-specific components from hidden representations. We evaluate this intervention across 10 open-source LLMs and 3 multilingual reasoning benchmarks. Results show consistent performance improvements, supporting our claim that disentangling language and reasoning enhances multilingual generalization.
@@ 1 Introduction
Recent advances in reasoning large language models—as exemplified by OpenAI’s o1/o3/o4 models (Jaech et al., 2024; OpenAI, 2025) and the DeepSeek-R1 series (Guo et al., 2025)—have significantly advanced the capacity of language models to handle complex reasoning tasks. These improvements stem from multi-stage post-training pipelines, notably supervised fine-tuning and large-scale reinforcement learning with reasoning-enhanced objectives (Guo et al., 2025; Kimi et al., 2025; Qwen, 2025; Muennighoff et al., 2025; Yeo et al., 2025). As a result, these models demonstrate increasingly human-like deliberative abilities and can generate extended chains of thought (CoT) to support long-horizon reasoning (Li et al., 2025a; Xu et al., 2025; Chen et al., 2025).
@@
Despite the remarkable advancements of LLMs’ reasoning capabilities, progress has been heavily concentrated in high-resource languages such as English and Chinese (Glazer et al., 2024; Phan et al., 2025). While these models exhibit strong performance in these dominant languages, they continue to face challenges in reasoning tasks across low- and mid-resource languages (Karim et al., 2025; Gao et al., 2025). This growing disparity in multilingual reasoning capabilities raises critical concerns: it hinders the global applicability of LLMs, exacerbates existing linguistic inequities in AI access, and perpetuates the marginalization of underrepresented languages due to limited data and investment (Okolo and Tano, 2024; Ghosh et al., 2025; Qin et al., 2025). Despite the significance of this issue, the multilingual reasoning gap remains largely under-examined in current research.
@@
In this work, we investigate the interplay between language and reasoning in LLMs, aiming to uncover the key factors that influence cross-lingual reasoning generalization. Our key hypothesis is that reasoning and language processing can be explicitly decoupled, allowing reasoning abilities—once acquired in a high-resource language—to transfer more broadly across languages. This idea is supported by cognitive neuroscience findings that the human brain’s language network—responsible for comprehension and production—remains largely inactive during reasoning tasks (Monti et al., 2009, 2012; Fedorenko et al., 2011; Amalric and Dehaene, 2019), and that human language itself is evolutionarily optimized for communication rather than for reasoning (Fedorenko et al., 2024).
@@
To further validate this hypothesis, we perform causal interventions within LLMs’ latent spaces during multilingual reasoning by subtracting language-specific components from each hidden state (§3). This intervention aims to disentangle linguistic features from the underlying reasoning process. Notably, as previewed in Figure 1, we observe consistent performance gains in multilingual reasoning across a diverse set of tasks, including mathematical problem solving (Shi et al., 2023), commonsense inference (Muennighoff et al., 2023), and knowledge-intensive question answering (Hendrycks et al., 2021). These improvements are robust across all 10 open-source LLMs, encompassing both reasoning-oriented and general-purpose LLMs. Moreover, the benefits generalize across 11 languages of varying resource availability. This consistent pattern provides compelling empirical evidence that reasoning and linguistic processing can be disentangled, enabling the effective transfer of reasoning capabilities from high-resource to low-resource languages.
@@
2 Decouple Language and Reasoning in the Activation Space
In §2.1, we identify language-specific subspaces by isolating components that consistently encode linguistic variation across inputs. In §2.2, we introduce a projection-based intervention that removes these components during inference to decouple language-specific information from the reasoning process. Finally, in §2.3, we provide empirical evidence that the removed components indeed correspond to language-specific signals, validating the effectiveness of our approach and the plausibility of language–reasoning decoupling in practice.
@@ 2.1 Language-Specific Subspace Identification
Building on prior works (Pires et al., 2019; Libovickỳ et al., 2020; Yang et al., 2021), the multilingual latent space $\boldsymbol{M}$ can be decomposed into two orthogonal components: (1) a language-agnostic subspace $\boldsymbol{M}_{a}$ that captures cross-lingual shared semantics, and (2) a language-specific subspace $\boldsymbol{M}_{s}$ that encodes language-dependent variations in linguistic expression. Following the formulation of Piratla et al. (2020); Xie et al. (2022); Zhao et al. (2024a), the decomposition objective is:
@@
2.2 Activation Ablation for Language-Reasoning Decoupling
Given that $\boldsymbol{M}_{s}$ characterizes the language-specific subspace within the model’s activation space, we leverage a ablation mechanism to decouple language-specific signals from multilingual reasoning. Specifically, for any hidden representation $\boldsymbol{h}$ derived from a multilingual input over a specific reasoning task, we project out its components along the subspace spanned by $\boldsymbol{M}_{s}$:
@@ 2.3 Verifying the Linguistic Nature of Removed Components
To assess the effectiveness of the projection-based intervention (Equation 3) in decoupling language-specific information from the model’s reasoning process, we conduct empirical validation from two complementary perspectives: (1) representation space visualization, and (2) language fidelity between input and output on a specific reasoning query. We perform this validation on Qwen-2.5-7B-Instruct, covering languages at different resource levels: French and Japanese (high-resource), Thai (medium-resource), and Swahili (low-resource).
@@
To further examine how this alignment in representation space affects the model’s behavior, we evaluate the language fidelity (Holtermann et al., 2024), a metric that quantifies consistency between the input and output languages. We use GlotLID (Kargaran et al., 2023), a multilingual language identifier supporting over 1,600 languages, to detect the language of model-generated responses. As shown in Figure 3, increasing the degree of ablation (i.e., projecting out more components of $\boldsymbol{M}_{s}$) leads to a substantial drop in language fidelity across all languages: models increasingly default to English in their responses, even when prompted in another language. This behavioral shift mirrors the representational alignment observed in Figure 2, confirming two key findings: (i) the removed components do encode language-specific signals, and (ii) once removed, the model tends to revert to its dominant pretraining language, i.e. English, as the default output anchor, consistent with prior work on multilingual representation aligment (Chen et al., 2023; Wu et al., 2024; Wang et al., 2024; Dumas et al., 2024; Wendler et al., 2024; Chen et al., 2024; Zhao et al., 2024b).
@@ 3 Causal Intervention within the Activation Space
We conduct intervention experiments across a diverse set of LLMs, covering both reasoning and non-reasoning models. Specifically, for non-reasoning (instruction-tuned) models, we include the Qwen-2.5-Instruct family (3B and 7B) (Yang et al., 2024) and the Qwen-3-Instruct series (1.7B, 4B, and 8B) (Team, 2025). For reasoning-oriented models, we study DeepSeek-R1-Distill (7B and 14B) (Guo et al., 2025), GLM-Z1-9B (GLM et al., 2024), and QwQ-32B (Qwen, 2025). Notably, for the Qwen-3 series, we employ the Thinking mode to enable explicit reasoning.
@@
We evaluate the impact of decoupling language-specific information from multilingual reasoning across three major benchmarks, including MGSM (Shi et al., 2023): mathematical reasoning, XWinograd (Muennighoff et al., 2023): commonsense reasoning, and M-MMLU (Hendrycks et al., 2021): knowledge-intensive question answering. We use Accuracy as the evaluation metric for all benchmarks. A detailed description of each benchmark is provided in Appendix E.
@@
Our intervention are implemented based on the vLLM framework (Kwon et al., 2023) for efficient inference. For each model, we ablate the language-specific components from mid-layer hidden states and then re-inject them at higher layers. This strategy preserves overall input-output consistency while allowing us to isolate the causal effect of language-specific information on reasoning. A detailed layer-wise analysis of this mechanism is provided in §4.2, and hyperparameter configurations can be found in Appendix F.
Table 1: Multilingual reasoning performance on MGSM datasets across different languages, before and after removing language-specific components within the activation spaces of the backbone models (+ Act. Ablation). The best results are highlighted in bold. The values in parentheses indicate language fidelity to indicate input-output consistency.
@@
+ Act. Ablation 91.6 85.2 79.2 79.2 83.2 75.2 79.2 78.8 41.2 64.8 8.8 69.67 (27.64%)
Qwen-3-4B-Thinking 95.6 88.8 79.2 83.2 88.0 81.2 85.6 85.2 72.4 0.688 32.0 79.56 (27.30%)
+ Act. Ablation 96.0 89.2 82.0 83.2 87.6 83.6 85.2 84.8 73.6 82.8 32.4 80.04 (27.31%)
@@
Language–reasoning decoupling consistently improves multilingual reasoning performance. This improvement holds across all model types and architectures. For non-reasoning models, both the Qwen-2.5 and Qwen-3 Instruct series exhibit clear gains in multilingual reasoning after language–reasoning decoupling. Similarly, for reasoning models, the intervention proves effective across models trained with different paradigms: both the DeepSeek-R1-Distill series, which are distilled from stronger models via supervised fine-tuning, and models optimized with large-scale reinforcement learning objectives like QwQ-32B, show substantial multilingual improvements.
Moreover, the benefits of language–reasoning decoupling are even more pronounced on XWinograd and M-MMLU benchmarks that emphasize commonsense inference and knowledge-intensive reasoning. Across these tasks, nearly all models show larger absolute gains in average accuracy after intervention. This suggests that linguistic interference is particularly detrimental in reasoning tasks requiring subtle context integration or factual grounding, and that removing language-specific noise in the representation space can yield stronger improvements under these conditions.
@@
Table 2: Multilingual reasoning performance on XWinograd and M-MMLU datasets across different languages, before and after removing language-specific components within the activation spaces of the backbone models (+ Act. Ablation). The best results are highlighted in bold. The values in parentheses indicate language fidelity to indicate input-output consistency.
@@ 4.1 Language-Specific Activation Negatively Correlates with Reasoning Accuracy
Figure 3: Effects of ablation strength on multilingual reasoning performance and output fidelity. Positive values indicate the proportion of language-specific components in $\boldsymbol{M}_{s}$ removed, while negative values indicate their injection (reinforcing language-specific signals). We report MGSM accuracy (reasoning performance, red), reasoning fidelity (green), and response fidelity (yellow).
Figure 4: Layer-wise impact of language–reasoning decoupling on MGSM accuracy and output fidelity. The x-axis denotes the starting layer index of the intervention. Most layers support effective decoupling that improves reasoning performance.
4.2 Layer-wise Effects of Language–Reasoning Decoupling
To investigate where language and reasoning are most effectively separable, we conduct a layer-wise analysis. Specifically, we apply the decoupling intervention to different layers of the model, from lower to middle to upper, and evaluate its impact on reasoning performance and output fidelity.
As shown in Figure 4, we observe that ablation applied at most layers, especially from lower to middle depths, consistently improves reasoning accuracy while maintaining stable output fidelity. This suggests that language–reasoning decoupling is broadly effective across the network and does not require precise targeting to specific layers to yield benefits.
@@ 4.2 Layer-wise Effects of Language–Reasoning Decoupling
These findings reaffirm the central claim that reasoning and language can be disentangled across a wide range of model depths, but highlight that middle layers offer the best trade-off: substantial gains in reasoning performance without sacrificing linguistic coherence. This also aligns with prior findings showing that multilingual models tend to form language-agnostic semantic spaces in their intermediate layers (Wu et al., 2024; Wang et al., 2024).

### v2

@@ Abstract
Multilingual reasoning remains a significant challenge for large language models (LLMs), with performance disproportionately favoring high-resource languages. Drawing inspiration from cognitive neuroscience, which suggests that human reasoning functions largely independently of language processing, we hypothesize that LLMs similarly encode reasoning and language as separable components that can be disentangled to enhance multilingual reasoning. To evaluate this, we perform a causal intervention by ablating language-specific representations at inference time. Experiments on 10 open-weight LLMs spanning 11 typologically diverse languages show that this language-specific ablation consistently boosts multilingual reasoning performance. Layer-wise analyses further confirm that language and reasoning representations can be effectively disentangled throughout the model, yielding improved multilingual reasoning capabilities, while preserving top-layer language features remains essential for maintaining linguistic fidelity. Compared to post-training methods such as supervised fine-tuning or reinforcement learning, our training-free language-reasoning disentanglement achieves comparable or superior results with minimal computational overhead. These findings shed light on the internal mechanisms underlying multilingual reasoning in LLMs and suggest a lightweight and interpretable strategy for improving cross-lingual generalization. Our code is available at: https://github.com/MuyuenLP/Language-Reasoning-Disentangle. Refer to caption Figure 1: Overview of our hypothesis and findings. Left: Motivated by cognitive neuroscience, we hypothesize that reasoning and language processing in LLMs can be disentangled. Right: To validate this hypothesis, we perform a causal intervention by removing language-specific components from hidden representations. We evaluate this intervention across 10 open-weight LLMs and 3 multilingual reasoning benchmarks. Results show consistent performance improvements, supporting our claim that disentangling language and reasoning enhances multilingual generalization.
@@ 1 Introduction
Recent advances in reasoning large language models—as exemplified by OpenAI’s o1/o3/o4 models (jaech2024openai; openai2025o3blog) and the DeepSeek-R1 series (guo2025deepseek)—have significantly advanced the capacity of language models to handle complex reasoning tasks. These improvements stem from multi-stage post-training pipelines, notably supervised fine-tuning and large-scale reinforcement learning with reasoning-enhanced objectives (guo2025deepseek; team2025kimi; qwen2025qwqblog; muennighoff2025s1; yeo2025demystifying). As a result, these models demonstrate increasingly human-like deliberative abilities and can generate extended chains of thought (CoT) to support long-horizon reasoning (li2025system; xu2025towards; chen2025towards).
@@
Despite the remarkable advancements of LLMs’ reasoning capabilities, progress has been heavily concentrated in high-resource languages such as English and Chinese (glazer2024frontiermath; phan2025humanity). While these models exhibit strong performance in these dominant languages, they continue to face challenges in reasoning tasks across low- and mid-resource languages (karim2025lost; gao2025could). This growing disparity in multilingual reasoning capabilities raises critical concerns: it hinders the global applicability of LLMs, exacerbates existing linguistic inequities in AI access, and perpetuates the marginalization of underrepresented languages due to limited data and investment (okolo2024closing; ghosh2025multilingual; qin2025survey). Despite the significance of this issue, the multilingual reasoning gap remains largely under-examined in current research.
@@
In this work, we investigate the interplay between language and reasoning in LLMs, aiming to uncover the key factors that influence cross-lingual reasoning generalization. Our key hypothesis is that reasoning and language processing can be explicitly disentangled, allowing reasoning abilities—once acquired in a high-resource language—to transfer more broadly across languages. This idea is supported by cognitive neuroscience findings that the human brain’s language network—responsible for comprehension and production—remains largely inactive during reasoning tasks (monti2009boundaries; monti2012thought; amalric2019distinct), and that human language itself is evolutionarily optimized for communication rather than for reasoning (fedorenko2024language).
@@
To further validate this hypothesis, we perform causal interventions within LLMs’ latent spaces during multilingual reasoning by subtracting language-specific components from each hidden state (§3). This intervention aims to disentangle linguistic features from the underlying reasoning process. Notably, as previewed in Figure 1, we observe consistent performance gains in multilingual reasoning across a diverse set of tasks, including mathematical problem solving (shi2023language), commonsense inference (muennighoff2023crosslingual), and knowledge-intensive question answering (hendrycks2021measuring). These improvements are robust across all 10 open-weight LLMs, encompassing both reasoning-oriented and general-purpose LLMs. Moreover, the benefits generalize across 11 languages of varying resource availability. This consistent pattern provides compelling empirical evidence that reasoning and linguistic processing can be disentangled, enabling the effective transfer of reasoning capabilities from high-resource to low-resource languages.
@@
2 Disentangle Language and Reasoning in the Activation Space
In §2.1, we identify language-specific subspaces by isolating components that consistently encode linguistic variation across inputs. In §2.2, we introduce a projection-based intervention that removes these components during inference to disentangle language-specific information from the reasoning process. Finally, in §2.3, we provide empirical evidence that the removed components indeed correspond to language-specific signals, validating the effectiveness of our approach and the plausibility of language–reasoning disentanglement in practice.
@@ 2.1 Language-Specific Subspace Identification
Building on prior works (pires-etal-2019-multilingual; libovicky2020language; yang2021simple), the multilingual latent space $\boldsymbol{M}$ can be decomposed into two orthogonal components: (1) a language-agnostic subspace $\boldsymbol{M}_{a}$ that captures cross-lingual shared semantics, and (2) a language-specific subspace $\boldsymbol{M}_{s}$ that encodes language-dependent variations in linguistic expression. Following the formulation of piratla2020efficient; xie2022discovering; liu2025principled, the decomposition objective is:
@@
2.2 Activation Ablation for Language-Reasoning Disentanglement
Given that $\boldsymbol{M}_{s}$ characterizes the language-specific subspace within the model’s activation space, we leverage a ablation mechanism to disentangle language-specific signals from multilingual reasoning. Specifically, for any hidden representation $\boldsymbol{h}$ derived from a multilingual input over a specific reasoning task, we project out its components along the subspace spanned by $\boldsymbol{M}_{s}$:
@@ 2.3 Verifying the Linguistic Nature of Removed Components
To assess the effectiveness of the projection-based intervention (Equation 3) in disentangling language-specific information from the model’s reasoning process, we conduct empirical validation from two complementary perspectives: (1) representation space visualization, and (2) language fidelity between input and output on a specific reasoning query. We perform this validation on Qwen-2.5-7B-Instruct, covering languages at different resource levels: French and Japanese (high-resource), Thai (medium-resource), and Swahili (low-resource).
@@
To further examine how this alignment in representation space affects the model’s behavior, we evaluate the language fidelity (holtermann2024evaluating), a metric that quantifies consistency between the input and output languages. We use GlotLID (kargaran2023glotlid), a multilingual language identifier supporting over 1,600 languages, to detect the language of model-generated responses. As shown in Figure 3, increasing the degree of ablation (i.e., projecting out more components of $\boldsymbol{M}_{s}$) leads to a substantial drop in language fidelity across all languages: models increasingly default to English in their responses, even when prompted in another language. This behavioral shift mirrors the representational alignment observed in Figure 2, confirming two key findings: (i) the removed components do encode language-specific signals, and (ii) once removed, the model tends to revert to its dominant pretraining language, i.e. English, as the default output anchor, consistent with prior work on multilingual representation aligment (chen2023bigger; wu2024semantic; wang2024sharing; dumas2024llamas; wendler2024llamas; chen2024journey; zhao2024large).
@@ 3 Causal Intervention within the Activation Space
We conduct intervention experiments across a diverse set of LLMs, covering both reasoning and non-reasoning models. Specifically, for non-reasoning (instruction-tuned) models, we include the Qwen-2.5-Instruct family (3B and 7B) (yang2024qwen2) and the Qwen-3-Instruct series (1.7B, 4B, and 8B) (qwen3). For reasoning-oriented models, we study DeepSeek-R1-Distill (7B and 14B) (guo2025deepseek), GLM-Z1-9B (glm2024chatglm), and QwQ-32B (qwen2025qwqblog). Notably, for the Qwen-3 series, we employ the Thinking mode to enable explicit reasoning.
@@
We evaluate the impact of language-reasoning disentanglement in multilingual reasoning across three major benchmarks, including MGSM (shi2023language): mathematical reasoning, XWinograd (muennighoff2023crosslingual): commonsense reasoning, and M-MMLU (hendrycks2021measuring): knowledge-intensive question answering. We use Accuracy as the evaluation metric for all benchmarks. A detailed description of each benchmark is provided in Appendix E.
@@
Our language-reasoning disentanglement are implemented based on the vLLM framework (kwon2023efficient) for efficient inference. For each model, we disentangle the language-specific components from mid-layer hidden states and then re-inject them at higher layers. This strategy preserves overall input-output consistency while allowing us to isolate the causal effect of language-specific information on reasoning. A detailed layer-wise analysis of this mechanism is provided in §4.2, and hyperparameter configurations can be found in Appendix F.
Table 1: Multilingual reasoning performance on MGSM datasets across different languages, before and after language-reasoning disentanglement within the activation spaces of the backbone models (+ L-R Disentangle). The best results are highlighted in bold. The values in parentheses indicate language fidelity to indicate input-output consistency.
@@
+ L-R Disentangle 91.6 85.2 79.2 79.2 83.2 75.2 79.2 78.8 41.2 64.8 8.8 69.67 (27.64%)
Qwen-3-4B-Thinking 95.6 88.8 79.2 83.2 88.0 81.2 85.6 85.2 72.4 68.8 32.0 79.56 (27.30%)
+ L-R Disentangle 96.0 89.2 82.0 83.2 87.6 83.6 85.2 84.8 73.6 82.8 32.4 80.04 (27.31%)
@@
Language–reasoning disentanglement consistently improves multilingual reasoning performance. This improvement holds across all model types and architectures. For non-reasoning models, both the Qwen-2.5 and Qwen-3 Instruct series exhibit clear gains in multilingual reasoning after language–reasoning disentanglement. Similarly, for reasoning models, the intervention proves effective across models trained with different paradigms: both the DeepSeek-R1-Distill series, which are distilled from stronger models via supervised fine-tuning, and models optimized with large-scale reinforcement learning objectives like QwQ-32B, show substantial multilingual improvements.
Moreover, the benefits of language–reasoning disentanglement are even more pronounced on XWinograd and M-MMLU benchmarks that emphasize commonsense inference and knowledge-intensive reasoning. Nearly all models show larger absolute gains in average accuracy after intervention. This suggests that linguistic interference is particularly detrimental in reasoning tasks requiring subtle context integration or factual grounding, and that removing language-specific noise in the representation space can yield stronger improvements under these conditions.
@@
Table 2: Multilingual reasoning performance on XWinograd and M-MMLU datasets across different languages, before and after language-reasoning disentanglement within the activation spaces of the backbone models (+ L-R Disentangle). The best results are highlighted in bold. The values in parentheses indicate language fidelity to indicate input-output consistency.
@@ 4.1 Language-Specific Activation Negatively Correlates with Reasoning Accuracy
Figure 3: Effects of ablation strength on multilingual reasoning performance and output fidelity. Positive values indicate the proportion of language-specific components in $\boldsymbol{M}_{s}$ removed, while negative values indicate their injection (reinforcing language-specific signals). We report MGSM accuracy (reasoning performance, red), reasoning fidelity (green), and response fidelity (yellow). Dashed lines indicate the original model’s performance without intervention.
Figure 4: Layer-wise impact of language–reasoning disentanglement on MGSM accuracy and output fidelity. The x-axis denotes the starting layer index of the intervention. Most layers support effective disentanglement that improves reasoning performance. Dashed lines indicate the original model’s performance without intervention.
4.2 Layer-wise Effects of Language–Reasoning Disentanglement
To investigate where language and reasoning are most effectively separable, we conduct a layer-wise analysis. Specifically, we apply the disentanglement intervention to different layers of the model, from lower to middle to upper, and evaluate its impact on reasoning performance and output fidelity.
As shown in Figure 4, we observe that ablation applied at most layers, especially from lower to middle depths, consistently improves reasoning accuracy while maintaining stable output fidelity. This suggests that language–reasoning disentanglement is broadly effective across the network and does not require precise targeting to specific layers to yield benefits.
@@ 4.2 Layer-wise Effects of Language–Reasoning Disentanglement
These findings reaffirm the central claim that reasoning and language can be disentangled across a wide range of model depths, but highlight that middle layers offer the best trade-off: substantial gains in reasoning performance without sacrificing linguistic coherence. This also aligns with prior findings showing that multilingual models tend to form language-agnostic semantic spaces in their intermediate layers (wu2024semantic; wang2024sharing).